The 8-year-old child with gastroenteritis is prescribed to receive 500 ml of lactated Ringer’s (LR) solution...

  1. The 8-year-old child with gastroenteritis is prescribed to receive 500 ml of lactated Ringer’s (LR) solution over the next 10 hours. How many milliliters per hour should the nurse administer?
  1. The HCP prescribed a dose of acetaminophen according to weight recommendations for the child weighing 48 lb. The package inserts state that the recommended dose is 15 mg per kg. What dose should the nurse administer?
Answer #1

ANSWER.

* Calculation of ml/hr.

          ml/hr = Volume(ml)/Time(hr).

    Here

        Volume(ml) = 500 ml.

         Time(hr) = 10 hr

As per above formula

            ml/hr = 500/10

                      = 50 ml/hr.

* Calculation of dose(mg) of drug required for a child.

         Weight of the the child = 48 lb

     Convert lb to kg for calculation

             1 lb = 0.454 kg

       So

           Weight of the child = 48 * 0.454

                                          = 21.8 kg

           Recommended dose = 15 mg/kg.

    Dose(mg) of drug required for this child = 15 * 21.8

                                                                    = 327 mg.